OpenStudy (anonymous):

A shipment of 12 microwave ovens contains 3 defective units. A vending company has ordered 4 of these units, and because all are packaged identically, the selection will be at random. What is the probability that at least 2 units are good?

OpenStudy (anonymous):

so usually 1/4 of the shipment is defective. so if we get 4 units. and 1/4 is bad. then we hav 4*1/4 is bad. which gives you a grand total of 1 bad one. so i think all the time you will have 2 good ones..

OpenStudy (perl):

no you re wrong

OpenStudy (lgbasallote):

are you classmates @perl & @EnchantixPixie ?

OpenStudy (anonymous):

No?

OpenStudy (lgbasallote):

:okay:

Directrix (directrix):

Probability of a defective oven = 3/12 = 1/4 = .25 Probability of a non-defective oven = .75 P( 2 Defects) = C(4,2) * (.25)^2 * (.75)^2 = .2109 P( 3 Defects) = C(4,3) * (.25)^3 * (.75)^1 = .0469 P(4 Defects) = C(4,4) * (.25)^4 * (.75)^0 = .0039 Probability of at least 2 defective ovens = .2109 + .0469 + .0039 = .2617